Determine the numerical value of the following expression without the use of a calculator: log10 (1000100) 100 + X 100 n=1 sin(πn) + 1 (−1)n ! · vuut 1000 Y m=1 1 cos(πm) 2 


1
Expert's answer
2021-09-14T06:04:18-0400


"(\frac{log_{10}(1000^{100})}{100}+\displaystyle\sum_{n-1}^{100}\frac{sin(πn)+1}{(-1)^n}).\sqrt{\displaystyle\pi_{m-1}^{1000}\frac{1}{cos(πm)^2}}......(1)"


solve

"\frac{log_{10}(1000^{100})}{100}=\frac{100 log_{10}^{1000}}{100}.....[log \space m^{n}=nlog\space m]\\\frac{100log_{10}^{10^{3}}}{100}\\=3log_{10}^{10}\\=3\space......[log_{10}^{10}=1]"


Now

"\displaystyle\sum_{n-1}^{100}\frac{sin(πn)+1}{(-1)^n})=\displaystyle\sum_{n-1}^{100}[\frac{1}{(-1)^n}]=0......(sin\space πn=0)" and "\sqrt{\displaystyle\pi_{m-1}^{1000}\frac{1}{cos(πm)^2}}=\displaystyle\pi_{m-1}^{1000}\frac{1}{cos(πm)}=1.......(cos\space πm=(-1)^m)"


putting these values in equation 1 we get

"(3+0)\times 1=3"

so

"(\frac{log_{10}(1000^{100})}{100}+\displaystyle\sum_{n-1}^{100}\frac{sin(πn)+1}{(-1)^n}).\sqrt{\displaystyle\pi_{m-1}^{1000}\frac{1}{cos(πm)^2}}=3"